# BanTH: A Multi-label Hate Speech Detection Dataset for Transliterated Bangla

### v1

@@ 2.6 Dataset Statistics
Following data validation, the dataset is split into train, test, and val datasets maintaining the standard train:test:val ratio. During the split, a stratification approach was followed to maintain a uniform distribution of hate and non-hate labels and further multi-label categories across all the splits. Further data statistics are given in Table 1.
@@ 3.3 Prompting Strategy
For the prompting of the LLMs, we adopt one of the following strategies: Why Positive/Why Negative Wang et al. (2023), HARE Yang et al. (2023), non-explanatory, explanation-based, CoT based or translation based. A comparative overview of the prompts are reported in Table 3. We consider GPT 4o and GPT-3.5 in prompting experiments. We design separate base prompts for binary and multi-label classification. Each of these base prompts is then extended to one of the prompting strategies. We consider both zero-shot and few-shot prompting techniques, E.1 and E.2. Overall, all the variations of the prompts can be categorized as follows:
Non-Explanatory. The non-explanatory prompting strategy forms the base-prompts for other prompting strategies, where we only ask the LLM to do binary or multi-label classification. The prompt also includes the hate speech definition, the target groups, and their definition in multi-label classification, key indicators of hate, and labeling instructions.
@@
HARE. In HARE-based prompting strategy we insert the line Let’s explain step by step, as adapted in Yang et al. (2023), along with the base-prompt and asked the LLM to return its reason for the corresponding binary or multi-label.
@@
Translation Based. LLMs are trained primarily on extensive formal text, which may limit their performance and explanatory capabilities when handling transliterated text. To address this issue, we propose a simple yet effective prompting strategy called the Translation-Based prompting strategy. This approach requires LLMs to first convert Bengali transliterated text into standard Bangla or English before proceeding with the usual classification. We test this proposed prompting method using Non-Explanatory, Explanatory, and CoT based prompting techniques.
@@ 4.2 Prompt Based LLM Performance
Our extensive experiments reveal that fine-tuned models outperform all prompt-based results for both binary and multi-label classification with a good margin (around 8% improvement than the prompting results). The few-shot prompting strategy enhances model performance compared to the zero-shot approach, although this is not true for all cases. Additionally, GPT-4 consistently outperforms GPT-3.5 across all experiments and prompting variations, yielding improvements of 8-10%. We also find that generating explanations can occasionally boost performance, but it may also lead to worse results in some instances. Our proposed prompting techniques achieve the best performance in zero-shot settings. Overall, our proposed TB-encoders demonstrate superior performance, surpassing all fine-tuned LMs and prompt-based results.
@@ 5 Error Analysis
Figure 4 illustrates per-label accuracy with the best performing model, TB-mBert. From the fig, it is evident that the classification accuracy for all the labels except Personal Offense label. Figure 3 indicates that most hate speech labels are categorized under Personal Offense. Consequently, the subset accuracy of the multi-label classification, as presented in Table 5, is lower than the per-label accuracy. This suggests that the model struggles to effectively classify samples of Personal Offense label.
@@
Figure 5 shows the result for video category-wise result on BanTH test dataset for both the binary and multi-label classification with the best performing model. As illustrated by the figure, the minimum accuracy for binary and multi-label classification is at least 80% and 50%, respectively, with Entertainment videos holding the highest accuracy in both the classifications, showing higher subset accuracy in multi-label classification (79.5%).
Table 6 shows the multi-label hate speech distribution in all three video categories. It is evident in the figure, that the number of Personal Offense in Entertainment video type is comparable to that of the other multi-labels of hate, whereas in the other two categories, News & Politics and People & Blogs, the number of Personal Offense hate is noticeably high. The overall improved accuracy in Entertainment category aligns with the result produced in Figure 4.
The results in Table 5 show that our Translation-based models, like GPT 4o Bangla/English Translation with explanations, perform nearly as well as top prompt-based models. As evidenced in Table 7 and Table 8, LLMs can naturally translate Bangla transliterated text into English, even without explicit instructions. However, when asked to translate, the LLM tends to give more literal translations. Table 8 also shows that poor translations often lead to incorrect classifications, especially with regional dialects or terms lacking direct English equivalents. English translations generally outperform Bangla ones. When not prompted to translate, the LLM focuses on identifying hate-related terms, which works for explicit hate speech but struggles with subtler forms of offensive content, leading to misclassification.
@@ Appendix A Additional Related Work
Appendix B BanTH Dataset Annotation Guidelines for Annotators
@@
Appendix C Details of BanTH Dataset
C.1 Target Group Description
The description of the target group with suitable example is give in the followings:
@@ C.1 Target Group Description
[Bangla Trans]: "Awamilig manei haramjada ci"
[Eng]: "The Awami League is a bastard."
@@
[Bangla Trans]: "Nastik koy kii"
[Eng]: "What does atheist say?"
@@
[Bangla Trans]: "Chele theke meye hoye jaiyen na abar. Mante parbo naa"
[Eng]: "Don’t turn from a boy into a girl. I can’t accept it."
@@
[Bangla Trans]: "Tui kothay salar bata"
[Eng]: "Where are you son of a bitch?"
@@
[Bangla Trans]: "Oder dhore dhore mere felo."
[Eng]: "Catch them and kill them"
@@
[Bangla Trans]: "Lojja nai Bangladesh er jonogoner"
[Eng]: "People of Bangladesh do not have shame"
@@
[Bangla Trans]: "Khanker pola Tor chahara dakatar moto chahara kano"
[Eng]: "Son of a bitch Why is your face like a robber?"
Note: [Bangla Trans] indicates the transliterated Bangla sentence and [Eng] is the corresponding English translation.
C.2 Video Category-wise Multi-label Distribution
@@ C.3 Wordcloud
(a) Political
(b) Religious
(c) Gender
(d) Personal Offense
(e) Abusive/Violence
(f) Origin
(g) Body Shaming
(a) English Political
(b) Religious(English Translated)
(c) Gender (English Translated)
(d) Personal Offense (English Translated)
(e) Abusive/Violence (English Translated)
(f) Origin (English Translated)
(g) Body Shaming (English Translated)
From Figure 6 and the corresponding English translations in Figure 7, we can visualize the mostly mentioned words for each label in the BanTH dataset. For Political label, people often referred to Hasina (ex-prime minister of Bangladesh), bnp (a major political party in Bangladesh), league/lig (wing of a political party in Bangladesh), dalal (broker), police (a law enforcement force), etc. As we have covered transliterated Bangla YouTube comments of July 2024, the BanTH dataset includes the public responses related to the political unrest of Bangladesh during that period. In Religious labeled sample texts, people often mention muslim (people following Islam), hindu (people following Hinduism), nastik (atheist), Allah (the one and only God in Islam), etc. Gender related hate comments are mostly related to female. In Personal Offense case, commenters address dalal (broker) and police (a law enforcement force) contemptuously using tui/tor. The same case is evident in Abusive/Violence labeled samples because the police showed violence during the student protest of July 2024. As transliterated Bangla is used by mainly Indian and Bangladeshi people, Origin related hate is often directed towards Bangladeshi and Indian people. For Body Shaming classified samples, we see people refer to others as mota (fat) and takla (bald) to body shame and it is directed mostly towards females as apu (sister) is one of the most used words. Although the BanTH dataset is concentrated on these types of hateful words, it may not be the case when the source is different.
@@ Appendix D Experimental Setup
For Language Models (LMs) baseline experiments, we fine-tuned a range of pre-trained LMs to establish baselines for binary and multi-label classification on the BanTH dataset. The models used include multilingual architectures like XLM-RoBERTaConneau et al. (2019) and mBERTDevlin et al. (2018), as well as language-specific models such as BanglaBERTBhattacharjee et al. (2022) and BanglishBERTBhattacharjee et al. (2022), which are designed for Bangla and code-mixed transliterated Bangla text, respectively. We also employed BanglaHateBERTJahan et al. (2022b), specifically optimized for detecting hate speech in Bangla, along with IndicBERTKakwani et al. (2020) and MuRILKhanuja et al. (2021), tailored for Indian languages. Additionally, we included models like CharBERTMa et al. (2020), and VAC-BERTBhattacharyya et al. (2023) to capture diverse model architectures, including lightweight models and character-level representations. The models were fine-tuned with a learning rate of 2e-5, using the AdamW optimizer. The maximum sequence length was set to 512 tokens, and training was performed for 5 epochs with a batch_size of 32. Early stopping was applied based on the validation loss to prevent overfitting. For prompt-based experiments, we used OpenAI API to get LLM responses for GPT-3.5 and GPT-4o models. For measuring the performance, we consider Macro-F1 & Accuracy metrics for binary classification and Macro-F1, Subset-Accuracy & Hamming Loss for multilabel classification. In our further pretraining experiments, we randomly masked 15% of the tokens and used Masked Language Model loss (MLM loss) as the pretraining objectives. We used learning rate of 1e-5, with batch_size = 32 and we trained the the TB-Encoder models for 5 epoch on the whole training dataset.
@@ Appendix E Prompts for Classification on BanTH dataset
E.1 Examples of Few Shot For Binary label Classification
HATE(true)
@@ 2. Amader team abar harlo, ki kharap khelse ora
E.2 Examples of Few Shot For Multi-label Classification
HATE(true)
@@
E.3 Base Prompt of Binary Classification
E.4 Base Prompt of Multilabel Classification
Base Prompt for Multiple Classification (0 shot) You are an expert at detecting and categorizing hate speech in Bangla text (written in Latin script/English letters). Your task is to classify texts into multiple categories of hate speech. HATE SPEECH CATEGORIES: 1. POLITICAL: Targeting individuals or groups based on political affiliation or beliefs 2. RELIGIOUS: Targeting individuals or groups based on religion (e.g., Hindu, Muslim, Buddhist, Christian) 3. GENDER: Targeting individuals or groups based on gender identity or expression 4. PERSONAL OFFENSE: Insults or attacks directed at specific individuals 5. ABUSIVE/VIOLENCE: Language that is abusive or incites violence 6. BODY SHAMING: Targeting individuals based on physical appearance or body type 7. ORIGIN: Targeting individuals or groups based on their place of origin, ethnicity, or nationality 8. MISC (Miscellaneous): Other forms of hate speech not covered by the above categories KEY INDICATORS OF HATE SPEECH: 1. Dehumanizing language or comparisons 2. Calls for violence or harm 3. Discriminatory slurs or epithets 4. Stereotyping entire groups negatively 5. Promoting supremacy of one group over others LABELING INSTRUCTIONS: 1. For each text, determine if it contains hate speech indicators for any of the defined categories 2. Assign (True) all relevant hate speech categories to the text 3. If no hate indicators are present, classify as non-hate speech (empty array for categories) SPECIAL CONSIDERATIONS: - Consider local context and cultural references - Be mindful of dialectal variations in spelling - Pay attention to code-mixing (Bangla-English hybrid phrases) - A single text may belong to multiple hate speech categories - Use the MISC category for hate speech that doesn’t fit into other categories You will receive an array of objects, each containing an ’id’ and ’text’. Analyze each text and provide the required classification.
E.5 CoT-Based Prompt
Additional Steps For CoT-Based Prompt CHAIN-OF-THOUGHT ANALYSIS STEPS: 1. Initial reading: Read the entire text and identify any potential hate speech elements. 2. Keyword identification: Highlight specific words or phrases that could indicate hate speech. 3. Context analysis: Consider the overall context and tone of the message. 4. Category mapping: Connect the identified elements to specific hate speech categories. 5. Multi-label consideration: Determine if the text fits into multiple categories. 6. Intensity assessment: Evaluate the severity or intensity of the hate speech, if present. 7. Final classification: Summarize the findings and assign relevant categories.
E.6 Translation-Based Prompt
@@
Additional Steps For Translation-Based Prompt You are an expert at detecting hate speech. You will be given an array of Bangla texts (written in Latin script/English letters). First, convert the Bangla Transliterated text into standard Bangla without losing the semantic meaning, then classify the text as either hate speech (true) or non-hate speech (false). Provide a reason for your classification and provide the translation.
Appendix F Additional Error Analysis
Transliterated Text English Translation Label Reason by GPT-4o+Exp Reason by GPT-4o+Exp+Translation Translation by GPT-4o+Exp+Translation
Corer jati sontras r osovbo jati. Ata kokono vlo hote parena A nation of thieves and terror is an uncivilized nation. This can never be good. Hate The text uses derogatory terms to describe a group as terrorists. The text uses dehumanizing language to describe an entire group as terrorists and uncivilized, which is a hateful stereotype. A nation of terrorists and uncivilized people. It can never be good.
Kothay gela khomota go Where has the power gone? Non-hate The text makes a vague statement without hate indicators. The text expresses a rhetorical question about power but does not express hatred or incite discrimination or violence against a specific group. Where did the power go?
Tomar punishment sure Your punishment is certain. Non-hate The text is a light-hearted comment without hate indicators. There is no expression of hate or incitement of violence; it seems more like light teasing or friendly banter. Your punishment is definite
Oboidho sontan Illegitimate child. Hate The text uses offensive language implying illegitimacy with a violent emoji. The term ’Oboidho sontan’ is an offensive term, which means ’illegitimate child’, used to insult. It is indicative of hate speech through personal offense. Illegitimate child
Vai ore niye naranari kortechen ken? Ore dekhleo to sick/abnormal dekha jay Brother, why are you fussing over him? He looks sick/abnormal just by seeing him. Hate The text calls someone sick/abnormal, which is body shaming. Text uses dehumanizing language calling someone sick/abnormal and is derogatory, indicative of body shaming or personal offense. Brother, why are you bothering with him? Just by looking at him, you can see he looks sick/abnormal.
Table 7: Table comparing the outputs generated by GPT-4o+Exp and GPT-4o+Exp+Translation[EN] both producing correct classifications
Transliterated Text English Translation Predicted Label by GPT-4o+Exp+Translation Reason by GPT-4o+Exp Reason by GPT-4o+Exp+Translation Translation by GPT-4o+Exp+Translation
Correct Output by GPT-4o+Exp+Translation[EN] and Incorrect Output by GPT-4o+Exp
@@ Appendix F Additional Error Analysis
world er shob deshei torun shomaj montri mp hoy khali Bangladesh chara eitar karon holo onar moto kichu political peraon jara sharajibon nijei kheye jete chay manusher jonno korte chay na In all the countries of the world, youths become social ministers and MPs, but except for Bangladesh. The reason behind this is a kind of political anarchy those who live their entire lives for themselves not for the welfare of people. Non-hate The text criticizes the entire political generation with derogatory implications. The text criticizes political leaders in Bangladesh for not focusing on youth leadership, but it does not incite hatred or use dehumanizing language. In all countries of the world, young people are made ministers or MPs, except in Bangladesh. The reason for this is that some political persons want to selfishly continue benefiting themselves instead of working for the people.
@@
Tomrao to bikri hoe geco You have been sold too Non-hate The text implies corruption and selling out, suggesting negative stereotyping. The phrase does not express hatred or discrimination towards any specific group or individual. They have also been sold out.
@@
Chup catukar khuni Shut up killer sycophant Hate The text seems to express frustration but lacks direct hate indicators. The phrase ’chup catukar khuni’ implies killing, which incites violence and hate. Shut up, murderous sycophant.
@@
Ore dhora hok ainer awtai ana hok Let him be caught and brought under justice Non-hate The text calls for arrest and punishment. The text talks about bringing someone under the law without expressing or inciting hate or violence. Let him be caught and brought under the law.
@@
Agey apnader mordhe indian gulare khuje ber koren First, find the Indians among yourselves Hate The text calls for identifying certain individuals without expressing hate. The text stereotypes and discriminates based on origin, which is a hate indicator. First, find the Indian infiltrators among yourselves.
Incorrect Output by GPT-4o+Exp+Translation[EN] and Incorrect Output by GPT-4o+Exp
@@
Myb eta India cal parar meye Maybe she is a girl from some slum in India. Hate The text derogatorily implies an Indian woman based on location, with a stereotype There is no evident expression of hate speech in identifying someone by a phrase or as a resident of a place. . Maybe she is a girl from India Calcutta.
@@
eto jhamelar ki orkar era beche thakle sob kiso vole abaro chori bodmashi korbe What’s the use of so much trouble? If they stay alive, they will forget everything and steal and do mischief again. Non-hate The text discusses obstacles without expressing hate. The text suggests that a group of people, if they remain alive, will engage in theft and misconduct again. It stereotypes and implies inherent criminal behavior. What is the point of all this trouble, if they stay alive, they will forget everything and steal and commit mischief again.
@@
Tora nijerai to political party er pasa chatis. You yourselves are the sycophants of political parties. Hate The text accuses others of opportunistic political affiliations. The text is a general statement about political affiliation without direct hate speech indicators. It appears accusatory but not incite violence or hatred. You people yourselves are involved with the political party.
@@
Bhai ja korso tai paiso..tmra joto marso oitar hisab k dibe ?? Brother, you got what you deserved… will you give an account for how much you have killed? Non-hate The text discusses the consequences of actions without hate indicators. This text contains implicit threat and justification for physical violence, possibly in a retaliatory context. Brother, you got what you did.. Who’s going to account for how much you hit?
@@
Na bnp na awami league Islami Andolon Bangladesh best. Neither BNP nor Awami League, Islami Andolon Bangladesh is the best. Non-hate The text expresses political preference without any hate indicators. This text promotes political supremacy, suggesting one political movement is superior to others. No BNP, no Awami League, Islami Andolon Bangladesh is the best.
Table 8: Table comparing the outputs generated by GPT-4o+Exp and GPT-4o+Exp+Translation[EN] both producing different classifications for a given Bengali Transliterated text.

### v2

@@ 2.6 Dataset Statistics
Following data validation, the dataset is split into train, test, and val datasets maintaining the standard 80:10:10 ratio. During the split, a stratification approach was followed to maintain a uniform distribution of hate and non-hate labels and further multi-label categories across all the splits. Further data statistics are given in Table 1.
@@ 3.3 Prompting Strategy
We adopt one of the following strategies: Why Positive/Why Negative Wang et al. (2023), HARE Yang et al. (2023), non-explanatory, explanation-based, CoT based or translation based. An overview of the prompts has been reported in Tab. 3. We consider GPT 4o and GPT-3.5 in the prompting experiments and have designed separate base prompts for binary and multi-label classification. Each of these base prompts is then extended to one of the prompting strategies. We consider both zero-shot and few-shot prompting techniques, reported in App. E. Overall, all the variations of the prompts can be categorized as follows:
Non-Explanatory. The non-explanatory prompting strategy forms the base prompts for other prompting strategies, where we only ask the LLM to do binary or multi-label classification. The prompt also includes the hate speech definition, the target groups, and their definition in multi-label classification, key indicators of hate, and labeling instructions.
@@
HARE. In HARE-based prompting strategy, we insert the line Let’s explain step by step, as adapted in Yang et al. (2023), along with the base-prompt and asked the LLM to return its reason for the corresponding binary or multi-label.
@@
Translation Based. LLMs are trained primarily on extensive formal text, which may limit their performance and explanatory capabilities when handling transliterated text. To address this issue, we propose a simple yet effective prompting strategy called the Translation-Based prompting strategy. This approach requires LLMs to first convert Bengali transliterated text into standard Bangla or English before proceeding with the usual classification. We test this proposed prompting method using Non-Explanatory, Explanatory, and CoT-based prompting techniques.
@@ 4.2 Prompt Based LLM Performance
Our extensive experiments reveal that fine-tuned models outperform all prompt-based results for both binary and multi-label classification with a good margin (~8% improvement over prompting). The few-shot prompting strategy enhances model performance compared to the zero-shot approach, although this is not true for all cases. Additionally, GPT-4 consistently outperforms GPT-3.5 across all experiments and prompting variations, yielding improvements of 8-10%. We also find that generating explanations can occasionally boost performance, but it may also lead to worse results in some instances. Our proposed prompting techniques achieve the best performance in zero-shot settings. Overall, our proposed TB encoders demonstrate superior performance, surpassing all fine-tuned LMs and prompt-based results.
@@ 5 Error Analysis
Figure 4 illustrates per-label accuracy with the best performing model, TB-mBert. From the fig, it is evident that the classification accuracy for all the labels except Personal Offense label. Figure 3 indicates that most hate speech labels are categorized under Personal Offense. Consequently, the subset accuracy of the multi-label classification, as presented in Table 5, is lower than the per-label accuracy. This suggests that the model struggles to effectively classify samples of Personal Offense label. As illustrated by Fig. 5, the minimum accuracy for binary and multi-label classification is at least 80% and 50%, respectively, with Entertainment videos holding the highest accuracy in both the classifications, showing higher subset accuracy in multi-label classification (79.5%).
@@
It is evident from Tab. 6 that the number of Personal Offense in Entertainment video type is comparable to that of the other multi-labels of hate, whereas in the other two categories, News & Politics and People & Blogs, the number of Personal Offense hate is noticeably high. The overall improved accuracy in Entertainment category aligns with the result produced in Fig. 4.
The results in Tab. 5 show that our Translation-based models, like GPT 4o Bangla/English Translation with explanations, perform nearly as well as top prompt-based models. As observed in Tab. 7, LLMs can naturally translate Bangla transliterated text into English, even without explicit instructions. However, when asked to translate, the LLM tends to give more literal translations. We also observe that poor translations often lead to incorrect classifications, especially with regional dialects or terms lacking direct English equivalents. English translations generally outperform Bangla ones. When not prompted to translate, the LLM focuses on identifying hate-related terms, which works for explicit hate speech but struggles with subtler forms of offensive content, leading to misclassification.
@@ Appendix A Additional Related Work
Appendix B BanTH Dataset Annotation Guidelines
@@
Appendix C Additional Details of BanTH
C.1 Formal Description of Target Groups
The description of the target group with suitable examples is given in the following:
@@ C.1 Formal Description of Target Groups
C.2 Category-wise Multi-label Distribution
@@ C.3 Wordcloud
From Figure 6 and the corresponding English translations in Figure 7, we can visualize the mostly mentioned words for each label in the BanTH dataset. For Political label, people often referred to Hasina (ex-prime minister of Bangladesh), bnp (a major political party in Bangladesh), league/lig (wing of a political party in Bangladesh), dalal (broker), police (a law enforcement force), etc. As we have covered transliterated Bangla YouTube comments of July 2024, the BanTH dataset includes the public responses related to the political unrest of Bangladesh during that period. In Religious labeled sample texts, people often mention Muslim, Hindu nastik (atheist), Allah (the one and only God in Islam), etc. Gender related hate comments are mostly related to females. In Personal Offense case, commenters address dalal (broker) and police (a law enforcement force) contemptuously using tui/tor. The same case is evident in Abusive/Violence labeled samples because the police showed violence during the student protest of July 2024. As transliterated Bangla is used by mainly Indian and Bangladeshi people, Origin related hate is often directed towards Bangladeshi and Indian people. For Body Shaming classified samples, we see people refer to others as mota (fat) and takla (bald) to body shame and it is directed mostly towards females as apu (sister) is one of the most used words. Although the BanTH dataset is concentrated on these types of hateful words, it may not be the case when the source is different.
@@ Appendix D Experimental Setup
For Language Models (LMs) baseline experiments, we fine-tuned a range of pre-trained LMs to establish baselines for binary and multi-label classification on the BanTH dataset. The models include multilingual architectures like XLM-RoBERTaConneau et al. (2019) and mBERTDevlin et al. (2018), as well as language-specific models such as BanglaBERTBhattacharjee et al. (2022) and BanglishBERTBhattacharjee et al. (2022), which are designed for Bangla and code-mixed transliterated Bangla text, respectively. We also employed BanglaHateBERTJahan et al. (2022b), specifically optimized for detecting hate speech in Bangla, along with IndicBERTKakwani et al. (2020) and MuRILKhanuja et al. (2021), tailored for Indian languages. Additionally, we included models like CharBERTMa et al. (2020), and VAC-BERTBhattacharyya et al. (2023) to capture diverse model architectures, including lightweight models and character-level representations.
The models were fine-tuned with a learning rate of 2e-5, using the Adam optimizer. The maximum sequence length was set to 512 tokens, and training was performed for 5 epochs with a batch size of 32. Early stopping was applied based on the validation loss to prevent overfitting. For prompt-based experiments, we used OpenAI API to get LLM responses for GPT-3.5 and GPT-4o models. For measuring the performance, we consider Macro-F1 & Accuracy metrics for binary classification and Macro-F1, Subset-Accuracy & Hamming Loss for multilabel classification. In our further pretraining experiments, we randomly masked 15% of the tokens and used Masked Language Model loss (MLM loss) as the pretraining objectives. We used a learning rate of 1e-5, with batch_size = 32 and we trained the the TB-Encoder models for 5 epochs on the whole training dataset.
@@ Appendix E Prompts for Classification on BanTH dataset
Base Prompt for Multilabel Classification (0 shot) You are an expert at detecting and categorizing hate speech in Bangla text (written in Latin script/English letters). Your task is to classify texts into multiple categories of hate speech. HATE SPEECH CATEGORIES: 1. POLITICAL: Targeting individuals or groups based on political affiliation or beliefs 2. RELIGIOUS: Targeting individuals or groups based on religion (e.g., Hindu, Muslim, Buddhist, Christian) 3. GENDER: Targeting individuals or groups based on gender identity or expression 4. PERSONAL OFFENSE: Insults or attacks directed at specific individuals 5. ABUSIVE/VIOLENCE: Language that is abusive or incites violence 6. BODY SHAMING: Targeting individuals based on physical appearance or body type 7. ORIGIN: Targeting individuals or groups based on their place of origin, ethnicity, or nationality 8. MISC (Miscellaneous): Other forms of hate speech not covered by the above categories KEY INDICATORS OF HATE SPEECH: 1. Dehumanizing language or comparisons 2. Calls for violence or harm 3. Discriminatory slurs or epithets 4. Stereotyping entire groups negatively 5. Promoting supremacy of one group over others LABELING INSTRUCTIONS: 1. For each text, determine if it contains hate speech indicators for any of the defined categories 2. Assign (True) all relevant hate speech categories to the text 3. If no hate indicators are present, classify as non-hate speech (empty array for categories) SPECIAL CONSIDERATIONS: - Consider local context and cultural references - Be mindful of dialectal variations in spelling - Pay attention to code-mixing (Bangla-English hybrid phrases) - A single text may belong to multiple hate speech categories - Use the MISC category for hate speech that doesn’t fit into other categories You will receive an array of objects, each containing an ’id’ and ’text’. Analyze each text and provide the required classification.
In-context Examples in Few-Shot Prompting for Binary Classification HATE(true)
@@ 1. Ei hindu gula ke desh theke ber kore dite hobe, era shobi desher shotru
Reason: Calls for the expulsion of religious groups (Hindus) and labels them all as enemies
@@ 2. Amader team abar harlo, ki kharap khelse ora
CoT-Based Prompt CHAIN-OF-THOUGHT ANALYSIS STEPS: 1. Initial reading: Read the entire text and identify any potential hate speech elements. 2. Keyword identification: Highlight specific words or phrases that could indicate hate speech. 3. Context analysis: Consider the overall context and tone of the message. 4. Category mapping: Connect the identified elements to specific hate speech categories. 5. Multi-label consideration: Determine if the text fits into multiple categories. 6. Intensity assessment: Evaluate the severity or intensity of the hate speech, if present. 7. Final classification: Summarize the findings and assign relevant categories.
In-context Examples in Few-Shot Prompting for Multi-label Classification HATE(true)
@@
Translation-Based Prompt You are an expert at detecting hate speech. You will be given an array of Bangla texts (written in Latin script/English letters). First, convert the Bangla Transliterated text into standard Bangla without losing the semantic meaning, then classify the text as either hate speech (true) or non-hate speech (false). Provide a reason for your classification and provide the translation.
Transliterated Text English Translation Pred. Reason by GPT-4o+E Reason by GPT-4o+E+T Translation by GPT-4o+E+T
@@
world er shob deshei torun shomaj montri mp hoy khali Bangladesh chara eitar karon holo onar moto kichu political peraon jara sharajibon nijei kheye jete chay manusher jonno korte chay na In all the countries of the world, youths become social ministers and MPs, but except for Bangladesh. The reason behind this is a kind of political anarchy those who live their entire lives for themselves not for the welfare of people. NH The text criticizes the entire political generation with derogatory implications. The text criticizes political leaders in Bangladesh for not focusing on youth leadership, but it does not incite hatred or use dehumanizing language. In all countries of the world, young people are made ministers or MPs, except in Bangladesh. The reason for this is that some political persons want to selfishly continue benefiting themselves instead of working for the people.
@@
Tomrao to bikri hoe geco You have been sold too NH The text implies corruption and selling out, suggesting negative stereotyping. The phrase does not express hatred or discrimination towards any specific group or individual. They have also been sold out.
@@
Chup catukar khuni Shut up killer sycophant H The text seems to express frustration but lacks direct hate indicators. The phrase ’chup catukar khuni’ implies killing, which incites violence and hate. Shut up, murderous sycophant.
@@
Ore dhora hok ainer awtai ana hok Let him be caught and brought under justice NH The text calls for arrest and punishment. The text talks about bringing someone under the law without expressing or inciting hate or violence. Let him be caught and brought under the law.
@@
Agey apnader mordhe indian gulare khuje ber koren First, find the Indians among yourselves H The text calls for identifying certain individuals without expressing hate. The text stereotypes and discriminates based on origin, which is a hate indicator. First, find the Indian infiltrators among yourselves.
@@
Myb eta India cal parar meye Maybe she is a girl from some slum in India. H The text derogatorily implies an Indian woman based on location, with a stereotype There is no evident expression of hate speech in identifying someone by a phrase or as a resident of a place. Maybe she is a girl from India Calcutta.
@@
eto jhamelar ki orkar era beche thakle sob kiso vole abaro chori bodmashi korbe What’s the use of so much trouble? If they stay alive, they will forget everything and steal and do mischief again. NH The text discusses obstacles without expressing hate. The text suggests that a group of people, if they remain alive, will engage in theft and misconduct again. It stereotypes and implies inherent criminal behavior. What is the point of all this trouble, if they stay alive, they will forget everything and steal and commit mischief again.
@@
Tora nijerai to political party er pasa chatis. You yourselves are the sycophants of political parties. H The text accuses others of opportunistic political affiliations. The text is a general statement about political affiliation without direct hate speech indicators. It appears accusatory but not incite violence or hatred. You people yourselves are involved with the political party.
@@
Bhai ja korso tai paiso..tmra joto marso oitar hisab k dibe ?? Brother, you got what you deserved… will you give an account for how much you have killed? NH The text discusses the consequences of actions without hate indicators. This text contains implicit threat and justification for physical violence, possibly in a retaliatory context. Brother, you got what you did.. Who’s going to account for how much you hit?
@@
Na bnp na awami league Islami Andolon Bangladesh best. Neither BNP nor Awami League, Islami Andolon Bangladesh is the best. NH The text expresses political preference without any hate indicators. This text promotes political supremacy, suggesting one political movement is superior to others. No BNP, no Awami League, Islami Andolon Bangladesh is the best.
Table 7: Comparison of predictions and reasoning generated by GPT-4o using Explanation (E) and Translation (T) prompting. Green indicates that the prediction (H:Hate and NH:Non-Hate) was correctly labeled by GPT-4o+E+T but incorrectly labeled by GPT-4o+E, and red indicates incorrect prediction by both the models.
(a) Political
(b) Religious
(c) Gender
(d) Personal Offense
(e) Abusive/Violence
(f) Origin
(g) Body Shaming
(a) English Political
(b) Religious(English Translated)
(c) Gender (English Translated)
(d) Personal Offense (English Translated)
(e) Abusive/Violence (English Translated)
(f) Origin (English Translated)
(g) Body Shaming (English Translated)